### v2

# Public-private funding models in open source software development: A case study on scikit-learn

Abstract
Governments are increasingly allocating funding for open source software (OSS) development to address concerns related to software security, digital sovereignty, and national competitiveness in science and innovation, amongst others. While announcements of governmental funding are generally well-received by OSS developers, we still have a limited understanding of OSS developers evaluate the relative benefits and drawbacks of such funding compared to other types of funding. This paper explores this question through a case study on scikit-learn, a Python library for machine learning, whose funding model combines research grants, commercial sponsorship, community donations, and a €32 million grant from the France’s artificial intelligence strategy. Through 25 interviews with scikit-learn’s maintainers and funders, this study makes two key contributions to research and practice. First, the study illustrates how the maintainers have weaved public and private funding into their project to ensure the continued provision of scikit-learn as a digital public good, as well as the importance of diversified funding and governance protocols for funding to safeguard the community ethos of the project. Second, it offers practical recommendations to various stakeholders. For OSS developer communities, it illustrates the benefits of a diversified funding model in balancing the merits and drawbacks of different funding sources. For companies, it serves as a reminder that sponsoring developers or OSS projects can significantly support OSS maintainers, who often struggle with limited resources and towering workloads. For governments, it emphasises the importance of funding the maintenance of existing OSS in addition to or exclusively funding the development of new OSS libraries or features. The paper concludes with suggestions for future research directions.
Keywords Open source software  $\cdot$ OSS funding models  $\cdot$ machine learning  $\cdot$ scikit-learn  $\cdot$ OSS sustainability
1 Introduction
Open source software (OSS) has moved up the priority list of governments in light of concerns about digital sovereignty [1], software security [2], and national competitiveness in science and innovation [3], amongst others. For example, the discovery of the Log4Shell vulnerability in November 2021, described as the “worst security problem in a generation” [4], prompted swift action from governments. In January 2022, the White House convened cross-sector stakeholders, including government agencies, Big Technology companies, and OSS foundations, to discuss the collective responsibility to improve the security of OSS, recognising its indispensable role as digital infrastructure whose maintenance largely depends on volunteers [5]. In June 2022, the newly formed European Working Team on the Digital Commons, comprising representatives from 19 member states of the European Union (EU) and the European Commission (EC), published its inaugural report, Towards a Sovereign Digital Infrastructure of Commons, highlighting the importance of funding OSS for bolstering the digital sovereignty of Europe [6]. Some months later, in October 2022, the German government established the Sovereign Tech Fund (STF), allocating a budget of €11.5 million for its first year alone to fund the development, improvement, and maintenance of open digital infrastructure to support digital sovereignty, software security, innovation, and digital democracy [7]. These developments underscore the growing recognition of OSS as digital public goods and the public sector’s responsibility to participate in funding its development [8].
The increasing interest and involvement of governments in funding OSS development has been generally well-received by OSS developers. However, the question of how OSS developers view the relative benefits and drawbacks of governmental funding compared to other sources of funding remains unexplored. This paper addresses this timely issue through a case study of scikit-learn, a popular Python library for machine learning (ML), which is described as “the Swiss army knife of ML” due to its widespread use in research and industry [9]. scikit-learn is developed by a team of core maintainers, who are primarily based at Inria, the French Institute for Research in Computer Science and Automation, and a global community of volunteers. Its sustainability as a community-led OSS project in the industry-dominated field of artificial intelligence (AI) [10] is in part attributed to its mixed funding model, including research grants, commercial sponsorship, and community donations, as well as the institutional backing of Inria since 2010, which have provided the means to employ full-time maintainers, to provide them with office space and IT resources, and to organise community events, amongst others. November 2021 was a turning point for scikit-learn. Recognising the significance of scikit-learn and OSS in general in AI research and development (R&D), the French government announced a €32 million grant in its AI strategy to support the expansion of scikit-learn into a comprehensive suite of open source tools for ML and AI for the benefit of researchers and developers in France and beyond [11].
Through 25 semi-structured interviews with the scikit-learn maintainers and their public and private funders, this paper explores the historical role of mixed public-private funding in supporting the scikit-learn project. The paper makes two key contributions with implications for research and practice. First, the study contributes novel findings about the design and implementation of a public-private funding model in an OSS project, which remains an unexplored topic in the literature. The analysis sheds light on the respective roles that public and private funders have played in supporting scikit-learn and how its maintainers have balanced and when necessary negotiated their funders’ diverse interests. Second, the study provides practical insights for practitioners. For OSS developer communities, it shows that a diversified funding model can act as an effective means for balancing the merits and drawbacks of different funding sources, as well as for containing the influence of funders, whether they are a company or a public body. For companies, it is a timely reminder that by sponsoring developers or an OSS project, they can make a substantial difference to OSS projects and maintainers, who struggle with towering workloads. For governments, it underscores the importance of funding maintenance of existing OSS and refraining from the novelty bias of funding new OSS libraries or features.
The paper has the following structure. First, it discusses prior work on the role of funding, amongst other mechanisms, in supporting the sustainability of OSS projects. Second, it presents the study design and the case of scikit-learn. Then, it presents the results and discusses their implications for research and practice. The paper concludes with a discussion of the key recommendations to practitioners and suggestions for future research on OSS funding models.
2 Related Work
2.1 The role of funding in OSS sustainability
The sustainability of OSS projects is a multi-faceted problem. With the increasing recognition of OSS as digital infrastructure [12, 13, 14], an extensive literature has examined the reasons for OSS project failures and successes. A range of factors can contribute to the sustainability of OSS projects [15], from the steady inflow of new contributors [16, 17, 18, 19, 20] to a number of maintainer “best practices,” including timely responses [21], mentoring [22, 20], and task recommendation [23]. Others point to the role that foundations play as “boundary organisations” [24], which enable collaboration between diverse contributors, including volunteers and for-profit companies, who collectively test, develop, and maintain OSS [25, 26, 27]. Conversely, prior work shows that a host of social, economic, and technological factors can cause OSS projects to fail (or “deprecate”), including usurpation by a competitor, a lack of time or interest amongst maintainers, conflicts amongst maintainers, or the obsolescence of the software, amongst others [28].
The financial health of an OSS project is an important variable for its sustainability, concerning if or to what extent its maintenance is funded or reliant on volunteers [29]. However, money has long been a contentious subject amongst OSS developers. On the one hand, many developers take pride in the origins of the OSS movement as a social movement [30], described as a “programmers’ paradise”, comprising “geeks”, “hackers”, and “hobbyists,” who do not work for money [31]. Many developers cite their political ideals [32], altruism [33], or their passion for working on hard problems [34] as intrinsic incentives for contributing to OSS. Many complain that “money ruins everything” [35]. On the other hand, the OSS ecosystem is awash with unfunded projects and overstretched volunteers, who struggle to keep up with a spectrum of “visible” and “invisible” tasks required to keep projects running [29, 36, 21].
The public disclosure of vulnerabilities in widely used OSS libraries, such as Heartbleed in April 2014 or Log4Shell in December 2021, which was described as the “worst security problem in a generation” [37], captured the attention of industry and governments alike. The latter brought to light the systemic issues of free-riding and under-investment in OSS as digital public infrastructure, and the need for cross-sector stakeholders, including the public and private sectors, to financially support the developer communities who build and maintain this infrastructure [8]. Indeed it has been estimated that OSS is present in 96% of codebases [38] and constitutes up to 90% of software stacks [39]. As Benjamin Bikinbine has argued, there is an urgent need “not just [for] investment in institutions, organisations, technologies, or innovations, but long-term and sustainable investment in the true source of their value: people” [40].
2.2 Funding models in OSS development
Funding models for OSS have been evolving in recent years, with individuals, public bodies, companies, and philanthropies playing a part in financially supporting OSS development and maintenance. Individuals—who may be OSS developers themselves or simply users of OSS—fund OSS developers by making micro-donations. For instance, the GitHub Sponsors scheme, launched in 2019, allows users to make one-off or recurring payments to individuals or organisations with sponsored profiles [41]. According to prior work, common incentives for donations via GitHub Sponsors include expressing appreciation, recognising a developer’s work, and encouraging future contributions [42]. By comparison, common incentives for maintainers to accept donations are to receive recognition or to provide a side income for their OSS work; while those that do not participate state that they do not need to be sponsored or they do not contribute to open source for money [42]. Other work shows that the amount that individuals donate is associated with the length of their use of an OSS [43]. The impact of micro-donations is still contested. On the one hand, they have been shown to shorten response times to issues [44, 45] and to increase maintenance-related activities [46]. On the other hand, the impact appears trivial [47] and donations tend to be ad-hoc, small, or legally complicated to receive [48].
The private sector has been the largest funder of OSS to date, funding OSS developers and projects in both direct and indirect ways. The impact of commercial investments in OSS are gauged to be significant; according to one estimate, companies in the EU invested around €1 billion in OSS development in 2018, which generated €65-95 billion for EU GDP that year [49]. Direct funding methods include sponsorship of OSS consortia or foundations, which operate on a fee-paying membership model [50]. Recently, companies have begun to establish FOSS Contributor Funds, which allow employees to nominate and vote on OSS projects that should receive funding [51]. By involving individual developers in the funding decision-making process, these funds facilitate a more democratic way for companies to support the OSS that they use in their products or services [52]. A shortcoming of these novel funds is that funding awards are typically limited to \$10,000 over a one-year time frame, which does not amount to adequate support for OSS projects in the long run. Alyssa Wright from Bloomberg’s FOSS Contributor Fund ([53]) has acknowledged this shortcoming, arguing that “true sustainability comes from a multi-sided commitment” and “while financial support is important, real engagement through contribution and dialogue is as crucial to the ecosystem’s survival as writing a check.”
Other ways companies support OSS projects is by sponsoring OSS developers or by allowing their employees to contribute to OSS at work [54, 55]. Sponsoring developers is a recognised strategy that companies employ to influence projects [56, 57] and to improve their reputation as OSS patrons, which in turn is considered to aid the recruitment of OSS developers [58, 59, 60]. In the mid-2000s, it was estimated that 40% of OSS contributors were paid [61]. We should expect this number to be much higher today, with companies like Google, Microsoft, and Amazon accounting for a rapidly growing number of OSS developers and contributions to OSS projects [62]. In addition, companies invest in projects through donations and joining project steering committees [63]. While OSS projects benefit from commercial support, maintainers seek to limit the influence or even the dominance of companies in projects [64, 65]. To manage the influence of companies, OSS developers make use of “boundary organisations,” such as consortia or foundations [50, 66]. These institutions provide the institutional infrastructure that limits the influence of a single company [67] and allows different interest groups, such as volunteers and for-profit companies, to collaborate together in OSS projects [24]. Such organisations enable OSS developers to “negotiate their relationship with companies and, when necessary, to defend their commons-based resources from unwanted influence” [40].
While the private sector has historically been the largest funder of OSS, the public sector has played a key role in funding OSS since the origins of the World Wide Web [68]. Governments support OSS through a range of channels, including research grants [69, 70] and dedicated funds, such as the Open Technology Fund in the USA, the EC’s Next Generation Initiative, or the STF in Germany [8]. Lately, OSS has moved up the priority list for governments due to concerns about software security [2] and digital sovereignty [1, 71]. For example, the EC’s OSS strategy argues that OSS can give “Europe a chance to create and maintain its own, independent digital approach and stay in control of its processes” [72] . In October 2022, the German government set up the STF to “invest in [the] digital commons, volunteer communities, and the open source ecosystem,” reasoning that thriving OSS ecosystems are key to digital sovereignty, securing software, economic growth, and digital democracy [7]. The same year, the European Working Team on Digital Commons, comprising 19 EU member states and the EC, underlined the potential of OSS to strengthen digital sovereignty by minimising vendor lock-in, increasing the bargaining power of states, and offering greater design control in the public sector [6].
The public sector’s embrace of OSS to build digital public services is a recognised strategy to stimulate competition in domestic software markets, which might otherwise be monopolised by industry giants, especially Big Technology companies [73]. For example, it has been estimated that the Circulaire 5608, a French law requiring government agencies to favour OSS when procuring software, led to an increase of nearly 600,000 OSS contributions from French developers per year as well as yearly increases in the national competitiveness of the French IT market [3]. Beyond governmental adoption of OSS, government investments in OSS are often targeted at stimulating national competitiveness. For example, in June 2023, President Emmanuel Macron of France announced a €40 million fund to create a “digital commons” to support the development of open source large language models to support domestic startups and to challenge the dominance of Big Technology companies in the AI industry [74].
While the recent surge in governmental interest in funding OSS has broadly been welcomed by practitioners, it remains an open question regarding how OSS developers view its relative benefits and drawbacks compared to other types of funding, as well as the impact thereof on OSS development and maintenance. As governmental interest and involvement in funding OSS development is currently on the rise, it is timely to examine this question and to offer practical insights from the perspective of OSS developers and maintainers. With this context in mind, this paper examines the role of public and private funding in the scikit-learn project, a widely used Python library for ML, and how its maintainers view the relative benefits and drawbacks of their mixed, public-private funding model.
3 Study Design
3.1 Research aims
This paper investigates the emerging role that governments are playing in funding OSS development and maintenance, and how OSS developers view the relative merits and drawbacks of governmental funding through a case study on the scikit-learn project. In software engineering research, a case study “is an empirical enquiry that draws on multiple sources of evidence to investigate one instance (or a small number of instances) of a contemporary software engineering phenomenon within its real-life context” [75]. Single case studies are suitable when one seeks to provide in-depth insights, generate new hypotheses, or develop theories about a critical, extreme, unusual, or revelatory phenomenon [76]. The scikit-learn project was selected as a revelatory case that could provide in-depth insights into the emergent role of the French government in supporting this OSS project and how its maintainers view the relative merits and drawbacks of the government’s funding compared to the other funding sources that have supported the project to date.
To this end, the study focuses on the following research questions:
  • RQ1: What are the interests of the public and private funders of the scikit-learn project, and how have they aligned or conflicted with the interests of the scikit-learn maintainers?
  • RQ2: How do the maintainers of the scikit-learn project view the relative benefits and drawbacks of public and private funding for OSS development?
  • RQ3: What lessons can be drawn from the scikit-learn project’s funding model for OSS practitioners?
3.2 Case presentation
The landscape of OSS in the field of ML/AI has evolved significantly in the last two decades. In 2007, a group of leading scientists lamented the dearth of OSS for ML and launched the Journal of Machine Learning Open Source Software to support OSS development in this scientific field [77]. Nowadays, OSS is considered indispensable to ML/AI R&D [78, 79] and, congruent with industry dominance in AI R&D [10], a significant number of OSS libraries for ML/AI have been initiated, hosted, and maintained by industry actors [80]. However, a noteworthy exception to this trend is scikit-learn, an OSS project that was born out of Inria, a public science body in France, and that is supported by a global community of volunteers. The following section provides an overview of the scikit-learn project.
scikit-learn is a Python library that implements algorithms for classification, regression, and clustering, as well as related tools for data preprocessing, model fitting, model evaluation, and visualisations. Initially called scikits.learn, the project was started by David Courapeau as a Google Summer of Code project in 2007. After a dormant period, it was relaunched as scikit-learn by Fabian Pedregosa, Gaël Varoquaux, Alexandre Gramfort, and Vincent Michel with the backing of Inria, with the first public release in February 2010. Nowadays, it is one of the most impactful and popular projects in the ML/AI OSS landscape. It is described as “the Swiss army knife of ML” due to its widespread use in research and industry [9]. By January 2022, it boasted a community of 2,250 contributors, 46,177 GitHub stars, and the most downloads (621,408,470) out of all Python libraries in this landscape between January 2017 and December 2021. By November 2023, the project boasted 2,752 contributors and 61,085 GitHub stars [81].
During the interviews, the maintainers explained that the sustainability of scikit-learn was not inevitable. There were many uncertainties throughout the project’s history, from the under-development of the ML OSS ecosystem in the early 2010s to market pressure in the later 2010s. One factor, amongst others, that contributed to the project’s sustainability was its funding model. The project has benefitted from several sources of funding [82]. Inria is the longest-standing funder, which has paid for the salaries of maintainers, office space, and events, amongst others, since 2010. In 2018, the scikit-learn consortium was established under Inria Foundation, comprising companies that pay an annual membership fee, in order to secure employment for maintainers, sustain the quality of software, and develop new features [83]. Since its founding, the consortium has included Dataiku, Microsoft, Nvidia, Intel, AXA, Boston Consulting Group, BNP Paribas Cardif, Hugging Face, Nvidia, Fujitsu, and Chanel. Other companies, such as the OSS consultancy Quansight Labs, have indirectly funded the project by sponsoring scikit-learn maintainers. The project also benefits from donations from community members via NUMFOCUS and student projects have been sponsored by the Google Summer of Code programme. In November 2021, the French government announced a €32 million grant in its AI strategy for the expansion of scikit-learn and the development of open source tools for data science [11]. The mixed funding of scikit-learn makes it a suitable case for a targeted analysis of the aforementioned research questions.
3.3 Data & analysis
3.3.1 Data collection
To answer the research questions, this study employs a mixed-methods approach, comprising document analysis and 25 semi-structured interviews with scikit-learn maintainers, commercial sponsors, and government officials between November 2021 and April 2023. Most interviews took place during on-site visits to the scikit-learn team’s office in Paris. The sampling approach sought to ensure complete coverage of the various stakeholder groups in the project—that is, the OSS developers (i.e. the scikit-learn maintainers), commercial sponsors (i.e. the consortium members), and public funders (i.e. Inria and the French government)—as well as diversity within each group. Interviews were conducted with maintainers with various tenures in the project, including some who had been involved for over 10 years and recent university graduates. One maintainer worked for a company that was a consortium member but spoke on his own behalf. From the consortium, interviews were conducted with three representatives from three companies with different sponsorship tiers, organisational sizes, and different sectors. In addition, interviews were conducted with two government officials who were responsible for the coordination of the AI strategy. The interviews were semi-structured, combining general questions about scikit-learn’s funding model and tailored questions for each interviewee [84]. To inform tailored questions for each maintainer, their contribution histories on GitHub’s Contributor Insights Page were reviewed; and for the funders, public information about their involvement in funding scikit-learn was reviewed. The author conducted and recorded every interview to aid the analysis and enhance the validity of the findings [76].
The interview data was supplemented and triangulated through detailed field notes from on-site visits and secondary documents to strengthen the validity of the findings [76]. The field notes stemmed from two on-site visits to the scikit-learn office in Paris-Saclay in March 2022 and 2023; two on-site visits to the offices of sponsors in September and December 2022; and attendance at the Global Partnership on AI Summit in November 2021, which was hosted by the French government and attended by a co-founder of scikit-learn. The secondary documents included both publicly available and non-disclosed documents, which interviewees shared concerning the French government’s AI Strategy and its funding package for scikit-learn. Additional information about scikit-learn and its funding model was collected from internet searches, included the scikit-learn website and company blogs, or provided by interviewees.
3.3.2 Qualitative data analysis
Throughout the 17 months of data collection, an iterative, integrated approach was employed to identify common themes in the interview data, field notes, and secondary documents [85]. Given the exploratory aims of this study, the integrated approach combined the application of grounded theory analysis [86] and a deductive, organising framework informed by prior literature about the different interests and involvement models of individuals, companies, and governments in the development of OSS [40, 58, 73, 24]. Thematic codes were continuously developed as data was collected, and in turn translated into themes until the point of saturation [87]. The resulting themes were member-checked to ensure their accuracy and resonance with research participants [88]. Throughout the research process, a social identity map was maintained by the author to exercise reflexivity regarding the potential influence of personal biases stemming from the author’s views and relationships built ith the scikit-learn team during the 17 months of data collection [89].
It should be noted that each stakeholder group was not assumed to be homogenous; instead, it was accepted that there could be divergent or even conflicting interests within each stakeholder group. For instance, the scikit-learn consortium comprises a range of companies, including startups and multinational corporations from different sectors. Prior work indicates that companies primarily pursue strategic interests by participating in OSS [90], and it was not assumed that the consortium members shared the same strategies in funding scikit-learn. Similarly, public funders, including Inria and government officials who were responsible for France’s AI strategy, were grouped together solely by their belonging to the French public sector, despite obvious differences between these stakeholders. Even the maintainers were a heterogeneous group, combining employees of Inria and external companies, each with their unique paths to and histories in the scikit-learn project. These theoretical groupings were used solely for analytical purposes and were iteratively refined during the 17 months of data collection to ensure their coherence and suitability for the analysis.
4 Results
The results are organised into three parts. The first part discusses the interests and stakes of the scikit-learn maintainers, while the second and third parts respectively report on the interests and involvement of the public and private funders. This structure sequentially responds to RQ1 concerning how the interests of the project’s funders align or conflict with those of the maintainers, and RQ2 concerning the relative merits and drawbacks of the different types of funding. With regards to RQ3, practical lessons from this analysis are presented in the subsequent Discussion section.
4.1 Stewards of the community: The incentives and stakes of the maintainers
The maintainers of scikit-learn are a heterogeneous group, comprising employees at Inria based at the Saclay science park on the outskirts of Paris and active community members, who, at the time of this research, were employed by companies, such as Nvidia, Quansight Labs, Microsoft, and Hugging Face. Several maintainers also held academic positions at universities and research labs. While some members of the core team worked on scikit-learn full-time, others balanced their roles as maintainers with their jobs. One maintainer explained that the maintainers, who are not employed by Inria, have varying levels of capacity to contribute because they have to balance competing demands on their time. Furthermore, the maintainers have had different journeys to scikit-learn. While some worked their way up to the role of maintainer as employees at Inria, others were invited based on their active involvement as community members. For example, Adrin Jalali explained that he began contributing in 2018 because he used scikit-learn as a data science consultant and got increasingly involved because “I got good feedback and there was a good community, so I started spending hours and hours and weekends on the project, and about six months later, I woke up to an email asking me if I wanted to be a core contributor.”
Maintainers shared the conviction that scikit-learn would not be what it is today without its community, which one maintainer argued was at the heart of the project’s DNA. During an on-site visit to the scikit-learn office, François Goupil, the community manager, showed me a dashboard of activity in the project, pointing to world maps of contributors. While the majority of the maintainers are based in Paris, France, the community contains contributors from all over the world (see Table 1). The maintainers lay great importance on this diversity and make efforts to cultivate the project’s identity as the collective effort of a global community. The dashboard also showed the involvement of various organisations, from research institutes and universities to startups and industry giants (see Table 2). Upon observing that the most active organisations are Chinese and US American Big Tech companies, I asked François if knowing how active these companies are had influenced his opinion about whether they contribute enough to the project. He responded frankly: while such companies contribute useful issues and pull requests, at the end of the day their contributions create more work for the already overloaded maintainers. It would be more helpful if they sponsored a developer or joined the consortium in order to provide more capacity to process the extra workload created by their contributions.
The maintenance of scikit-learn is a multifaceted job, spanning expertise in software engineering and statistical theory as well as social tasks like contributor mentorship and community animation, amongst others. François explained, “Of course, software development itself is the major part of the workload—going through issues, pull requests, and responding to community feedback—but lots of other work goes into the project. Community animation is a huge part. After all, it’s a community project.” He elaborated that in some ways the maintainers act like “editors” since they do not rush to implement issues or pull requests; rather, they carefully assess the needs and significance of issues and pull requests, liaise with maintainers of key dependencies, attend conferences and workshops, and read papers to stay up to date with research trends. “We don’t want to make easy, small fixes that crash in a few months,” he explained. Furthermore, sponsored developers from companies make focused contributions, such as the development of new features, which are collectively agreed upon by the maintainers and the sponsoring company. For example, at the time of this research, Nvidia sponsord a full-time maintainer who was building a plug-in to enable alternative computational backends for scikit-learn to enable compatibility with Nvidia’s AI hardware accelerators.
The demands of the job begs the question of incentives. Some maintainers explained that the work resonated with their values. The first time I met Gaël Varoquaux, he described scikit-learn as a political project that aims to provide a public alternative to the tools and products offered by industry giants. He stressed that scikit-learn was not the fruit of commercial interest and that decision-making about the project must always “come from the community.” The community ethos was underlined by other maintainers. Julien Jerphanion explained that only a minority of the scikit-learn community are paid for their contributions. “Even some of the maintainers are working on it in their spare time. The major part of the labour is based on benevolence of people working in their free time and not asking to get paid.” Community members perform many non-technical tasks, such as creating onboarding resources, helping the triaging team, and organising events. Many also make financial contributions in the form of micro-donations via NUMFOCUS. While “the major stuff is not funded through NUMFOCUS”, one maintainer explained that it has been useful in funding marketing, events, and so far one internship for underrepresented groups.
Beyond scikit-learn itself, several maintainers explained they care deeply about contributing upstream to the wider scientific Python ecosystem. During an on-site visit in September 2022, the maintainers had just returned from EuroSciPy, an annual gathering of European stakeholders in the scientific Python community. One maintainer spoke fondly of “putting faces to GitHub IDs.” Several maintainers mentioned that they felt a responsibility to support their dependencies, such as numpy and scipy, and that they collaborate with these projects in numerous ways beyond the software itself, including by applying for grants together. In addition, some of the maintainers had initiated “satellite projects,” such as joblib and skrub. However, their ability to support the ecosystem is constrained by their limited resources and hefty workload. Julien explained that while projects like scikit-learn, numpy, and scipy are very important, “There are many which are not visible to the public, for example OpenBLAS which is fundamental and only maintained by two people, who have very specific expertise. This is tremendous software work, but it is invisible to the public. Projects like this are cornerstones for many projects. They are the foundations; they are, like, infrastructure — you do not see them if there is no problem in your daily life.”
The maintainers also have personal interests in maintaining scikit-learn. Several maintainers explained they became maintainers because of their scientific interest in ML, and they like the scientific culture of the project and at Inria. Many of the maintainers have academic backgrounds, some work in research labs, and some teach at universities. One maintainer explained that while he may not become a millionaire contributing to open source, he had a comfortable life and enjoyed the challenge of the job, whilst another maintainer said that he enjoyed working at the intersection of statistical theory and software engineering, which was lacking in his previous job in industry. For many, being a maintainer of scikit-learn is also a source of pride, given the project’s reputation and impact in the field of ML/AI.
The diversity of incentives in the community is understood by the maintainers as a strength of the project, where decision-making is made based on the consensus of a diverse community. Adrin explained, “Even as a core developer, I don’t have enough credit to just be like, ‘Hey, we should do this.’ It’s a slow project with hundreds of 1000s of users. You can’t just change things.” Furthermore, the active involvement of the community was stressed as a key factor for the project’s sustainability. For example, Julien explained, “It is important to have people get paid for the work they do for scikit-learn. But if there’s only people that are getting paid, then something’s missing. I think what’s important it’s that there’s an understanding that an open source project is a commons and people can get involved if they want to.”
Table 1: Most active countries in the scikit-learn community [81]
Stargazer Issue Creator Pull Request Creator
1
China (23.7%)
USA (33.7%)
USA (34.0%)
2
USA (21.8%)
Germany (7.5%)
India (8.9%)
3
India (7.9%)
UK (7.0%)
Germany (8.5%)
4
Germany (4.4%)
India (6.8%)
France (6.9%)
5
Brazil (3.8%)
France (5.5%)
UK (6.4%)
6
UK (3.7%)
China (5.4%)
Canada (4.5%)
7
Canada (3.3%)
Canada (4.5%)
China (2.6%)
8
France (3.2%)
Netherlands (2.5%)
Japan (2.5%)
9
Japan (2.5%)
Switzerland (2.2%)
Switzerland (2.3%)
10
Russia (1.3%)
Australia (2.0%)
Netherlands (2.2%)
Table 2: Most active countries in the scikit-learn community [81]
Stargazer Issue Creator Pull Request Creator
1
Google (1.5%)
Google (2.5%)
Google (3.7%)
2
Microsoft (1.5%)
Microsoft (2.0%)
Microsoft (2.5%)
3
Tencent (0.9%)
Amazon (0.7%)
Inria (0.9%)
4
Alibaba (0.8%)
DeepMind (0.7%)
Meta (0.8%)
5
Amazon (0.8%)
IBM (0.6%)
Amazon (0.8%)
6
ByteDance (0.8%)
Johns Hopkins (0.6%)
Dataiku (0.7%)
7
Baidu (0.6%)
Tencent (0.5%)
UC Berkeley (0.6%)
8
Meta (0.5%)
ETH Zurich (0.5%)
Shopify (0.6%)
9
AWS (0.4%)
Uni. of Washington (0.4%)
MIT (0.6%)
10
Tsinghua University (0.4%)
NeuroData (0.4%)
AWS (0.6%)
4.2 From Inria to France’s AI strategy: The role of public funding in scikit-learn
The public sector has funded the project since its very beginning. While it initially started as scikits.learn during a Google Summer of Code project in 2007, it was abandoned and later exhumed by research fellows at Inria in 2010. To date, Inria has supported scikit-learn by employing staff, providing office space, and sponsoring events. A rough estimate of their sponsorship since the beginning is €1.5 million. “So, it was initially with public money from Inria from French public research grants that scikit-learn was developed,” explained François. Olivier Grisel, a maintainer, similarly explained that the public sector has played a long role in funding scikit-learn through Inria: “This is not new with the AI strategy.” Inria’s support role is not limited to funding. Through the Inria Foundation, which was established in 2017 to mobilise financial resources to support ambitious research projects, it has provided the institutional infrastructure for the scikit-learn consortium and its commercial sponsorship since 2018. One maintainer explained that while Inria may not be widely known in the community, its historical role in supporting scikit-learn cannot be overlooked.
The French government’s announcement that it would invest €32 million in the expansion of scikit-learn marked a new phase in the public sector’s involvement in the project. The grant is part of the strategy’s aim to cement France’s national competitiveness in AI and to increase AI adoption in the French economy. The strategy highlights that scikit-learn is an asset, which is used by millions of data scientists on a daily basis and which has been supported by Inria since its beginning. A government official explained that the funding for scikit-learn and related open source tools was intended to increase the adoption of AI throughout the country. By virtue of open source, it was acknowledged that researchers and businesses beyond France would benefit, too. In addition to economic goals, the government official explained that digital sovereignty was a key consideration for funding scikit-learn and the digital commons more widely. This was underscored by secondary documents that interviewees shared with me that contextualised the government’s funding for scikit-learn as part of a broader strategy to develop sovereign digital commons in France and Europe. The same government official explained that there is a “real challenge” with the dependence on technologies provided by US Big Tech companies, even including open source frameworks like PyTorch and TensorFlow, in the field of AI and by investing in the digital commons the government aspired to address such dependencies.
Another government official explained the funding decision was informed by a multi-stakeholder process involving stakeholders from research and industry. “It was not just the government. It was a dialogue between several actors—with the AI strategy team, the Confiance IA project, [^1] Confiance IA is a government-sponsored programme to develop and industrialise trustworthy AI technologies. Inria, with scikit-learn, and so on.” However, the coordinators of the AI strategy faced hurdles within the government to get support for the grant: “Financing open source means financing a product that cannot be valued on the market and a priori it is not clear what its success will be. So, it’s very problematic in the sense that if the government finances it in a substantial way—here we are talking about €32 million—the government wants to know where the money is going and to achieve a measurable return on investment.” It took months of coordination and negotiation to eventually get the sign-off from the finance ministry. A maintainer explained that the sum would in fact be double of the announced sum, given a requirement for the scikit-learn team to double the sum, with one half from the private sector and the other half from other public bodies in Europe.
Several maintainers mentioned that they were pleased to receive support from the government. Julien commented, “scikit-learn is a global project, but it started at Inria, so it’s nice to get some recognition from the government. It’s a good signal that the French state is getting involved in funding projects like this.” François commended the government’s requirement to double the funding with equal halves stemming from public and private sources across the EU. “I think this is a good vision because we are a community with different interests. I think it would be dangerous for us to be exclusively funded by the private sector or to be exclusively funded by the French government, because we have many good contributors who are not French, and they could have the feeling that, ‘Okay, now it’s a French government project, they are buying it.”’ Meanwhile, Adrin highlighted that the funding had secured the sustainability of the project for at least five years, which meant that they could make long-term plans regarding recruitment and project development.
However, the government’s funding of scikit-learn has not been without its challenges. Some of the maintainers commented that the government’s focus on high-level public policy goals meant that some of their demands were removed from their daily practices or expertise as OSS developers. Initially, the government officials had neglected maintenance and requested that the scikit-learn team expand the library to deep learning (DL) to provide a competitive alternative to Meta’s PyTorch [^2] N.B. Meta donated PyTorch to the Linux Foundation in September 2022. and Google’s TensorFlow. The scikit-learn team pushed back against this request, reasoning that they should focus on their strength in ML and consolidate their maintenance capacity. The government was receptive to this feedback. A government official told me, “We acknowledge that scikit-learn is specialised in ML, and it should not compete against TensorFlow and PyTorch in DL.” Moreover, one maintainer mentioned that the operationalisation of government objectives “is not easy to do, and sometimes it really does not make sense.” For example, on the one hand, the government wants the scikit-learn team to develop open source data science tools, while on the other hand it requires that scikit-learn does not compete with French companies. In practice, this is not that easy to do because the tools that they will build will be open source alternatives to proprietary products and services offered by French companies. It was a tricky demand to balance these two expectations. Another challenge for the maintainers was the pace of delivery. One maintainer expressed that even 17 months after the initial announcement, they still had not received the first instalment of the grant. While scikit-learn can support itself through its mixed funding model, he explained that the slow pace of funding could be problematic for other projects in more precarious financial situations.
While the interviewed sponsors saw the government’s funding positively, two representatives shared their concerns about the potential politicisation of the project owing to the French government’s substantial support for the project. One representative underlined the need to balance the French government’s concern for digital sovereignty with the norms of the global community, while another argued that imposing policy goals on the OSS project would not be received well by the community. These concerns were acknowledged by maintainers who I discussed these concerns with, who insisted that the community ethos of the project remained at the heart of their discussions with the government.
4.3 The muscle of industry: The role of the scikit-learn consortium
The scikit-learn project benefits from direct and indirect funding from companies that use scikit-learn in their business products and services. Some companies, such as Nvidia and Quansight Labs, have sponsored developers to contribute to scikit-learn. However, the most substantial funding has come from the scikit-learn consortium, which operates on an annual fee-based membership. François explained that its creation in 2018 addressed two pain points that the project faced: first, it would provide more stable funding for the project beyond individual research grants; and second, it would enable the project to offer permanent contracts to the maintainers, which was not possible previously at Inria. He explained the consortium offers companies “a way of giving back to the community and behaving responsibly towards a software tool that is part of their core business or their products.” In addition to financial support, the through the consortium’s Advisory Committee and the Technical Committee, companies advise on the project’s technical roadmap and share their industry use cases with the maintainers.
Companies pursue a combination of goals by joining the scikit-learn consortium. In both public blogs and interviews, companies cited reciprocity as a key motivation. Leo Dreyfus-Schmidt, the Dataiku representative, told me, “From our perspective, we think it’s a great opportunity to give back to the community because we are very much aware that we rely on open tools, especially scikit-learn.” A representative from another company added, “Companies pay millions in commercial licences, so why not support the OSS projects we use?” Companies also seek to capitalise on the reputational benefits of supporting scikit-learn, which in turn helps with hiring developers, who want to work at a company where OSS is considered important. François explained, “A lot of people come from this open source background and if you see that a company is promoting such an open source project, you might say, ‘Okay, maybe, I would be happy working with them. They could be okay with me contributing to OSS at work.”’ This incentive was viewd by some maintainers as more important for companies that did not already have a strong reputation in ML, since it would make them more credible and attractive to jobseekers.
Since all the companies in the consortium are users of scikit-learn, a key incentive for funding the project is to sustain its maintenance since it is a key dependency for their business. Mike McCarty from Nvidia shared that his concern about unmaintained dependencies “keeps me up at night” and cited NumPy as an example of a critical library that had been under-maintained for some time. In addition to maintenance, joining the consortium opens the door for a direct exchange between companies and the scikit-learn maintainers about industry use cases. Julien explained that he appreciates this exchange because “it shows culturally that those companies are not just using the software because it’s free, and it works; they are giving money to make the project sustainable.” Similarly, François commented, “[their] feedback is really rooted in something real. They have needs, they are working with customers on real projects, they need it to work.” Leo Dreyfus-Schmidt from Dataiku suggested, “I think that it’s a win-win for everybody here because putting myself in scikit-learn’s shoes, they don’t only get money, but they also get the perspective of people in industry, which for scikit-learn is a very different type of feedback from the one they get from the open source community. And it’s a win for all the consortium partners because it acts as a way to bring people together, which otherwise I wouldn’t know how that would happen.” Similarly, another representative explained, “It is good for the scikit-learn team to hear feedback from industry; not just small issues but how their tools are being used and what real world challenges are.” Several functionalities have been built into the scikit-learn library as a result of these exchanges.
Some companies have explicitly sought to influence the technical roadmap. For example, Nvidia sponsors a full-time maintainer who is building a plug-in to enable alternative computational backends. Mike McCarty explained, “One thing that Nvidia is very interested in is scikit-learn’s effort to accelerate with GPUs. I would love to see at the end of the day scikit-learn to be accelerated natively with a very nice user experience, kind of like what we have with PyTorch, which is the canonical example of a library that you can install and if you happen to have a GPU, you can accelerate it without changing your code.” Guillaume Lemaitre, a maintainer, described receiving funding and a maintainer at the same time as “a win-win situation.” However, sponsors’ demands are not always aligned with the norms of the project. Two maintainers explained that in the past, a sponsor had requested modifications to resources that amounted, in their view, to free advertising for the company, which was not acceptable for the maintainers. While disputes of this kind have been rare, one maintainer mentioned that they can lead to the termination of their sponsorships and one company left the consortium following a disagreement.
In order to safeguard community interests, the maintainers have designed governance protocols to limit the powers of the consortium. Adrin explained, “We don’t want the technical committee to have power over the core contributor team as long as [the maintainers] have a consensus. If there’s two thirds majority in cast vote, for whatever vote we do, then the technical committee doesn’t step in…The technical committee doesn’t do much in the project because the power is given to the [maintainer] team and that’s by design.” These governance measures are important to maintain the trust of the community and to mitigate fears that the project had been bought out. François explained that as a community-led OSS project, they communicate to consortium members that their feedback is valued but they cannot guarantee that they will implement it. For example, the maintainers organised a workshop about MLOps since this was a topic that the consortium members had expressed an interest in. “MLOps will probably end up in the technical roadmap in some way; we will say, ‘Okay, we can do this, we can do that.’ But it’s a discussion and in the end the team will decide if it’s worth it or not,” he explained. Through these governance measures, the maintainers ensure that the project remains community-driven while still considering and incorporating valuable feedback from its commercial sponsors.
5 Discussion
5.1 Implications for Research
This case study on scikit-learn illustrates the role that funding can play in supporting the sustainability and success of a non-commercial, community-led OSS project in industry-dominated fields like AI [10].
5.1.1 The design and role of public-private OSS funding models
This study makes novel contributions to the academic literature on OSS funding models. In particular, it contributes novel findings on the joint role of the public and private sector in funding the development and maintenance of an OSS project, a hitherto unexplored funding model. The analysis sheds light on the diversity of interests that public and private funders pursue in funding OSS development and maintenance, and their degree of alignment with OSS developers and maintainers. Extending prior work on incentives in OSS development, the case study illustrates that while the maintainers are primarily motivated by intrinsic incentives, such as their personal interest and values [91], companies primarily pursue economic and technological goals [58], such as maintenance and feature development, whilst governments are concerned with political and economic policy goals, including digital sovereignty [1], national competitiveness [3], and economic growth [73]. While the stakeholders pursue different goals, they are not necessarily in conflict, as they share the goal of sustaining scikit-learn. For example, while the maintainers and the companies do not share the government’s policy goals or digital sovereignty concerns, its long-term investment in scikit-learn as a digital public good aligns with the maintainers’ and companies’ interest in its sustainability for research and innovation purposes. Similarly, while the maintainers and the government do not share the companies’ business interests, the project benefits from companies’ feedback and funding for maintaining scikit-learn and expanding its functionalities.
In addition, it highlights the integral role played by the Inria Foundation as the institutional home of the scikit-learn consortium and the scikit-learn project’s governance protocols, which have served as crucial mechanisms for the maintainers to contain the influence of funders and uphold the project’s community ethos. This finding extends prior work on the role of foundations as “boundary organisations” that enable collaboration between OSS developers and for-profit companies with new evidence of the role of these institutions in handling the politics of money as an instrument of influence in community-led OSS projects [50, 24]. Moreover, the diversified funding model not only ensures financial stability for the project but also acts as a safeguard when disagreements arise between maintainers and funders. This diversification protects the project from the risk of over-reliance on the support of a single company.
5.1.2 Research design for future research on OSS funding
There is a dearth of research on OSS funding, relative to other topics in the extensive literature on the economics of OSS [92]. With the emergence of new funding models, from FOSS Contributor Funds to governmental funds like the STF, it is timely to empirically investigate OSS funding from multiple angles, including funders, incentives, funding models, and impact(s), amongst others. To this end, this paper contributes novel insights and a research design that can inform future studies on OSS funding. As discussed below, while the uniqueness of scikit-learn limits the generalisability of the case-specific findings, there are aspects of the scikit-learn experience that will resonate with OSS developers in general, such as the processes of balancing the varied interests of funders while preserving community values.
5.2 Implications for Practice
5.2.1 Implications for OSS developer communities
The experience of scikit-learn shows that a diversified funding model can act as a useful mechanism for safeguarding the project’s independence and community-driven ethos, whilst benefiting from the financial support of public and private funding, which have their relative merits. On the one hand, companies focus on specific technical problems related to maintenance and the development of new functionalities. As one maintainer explained, as users of scikit-learn, they need it to work. Maintainers shared their appreciation for the exchanges with commercial sponsors, which are “rooted in something real.” However, in some cases, companies make requests that the maintainers consider to be at odds with the interests of the project and refuse to implement, which can create tension and uncertainty over whether the company will sponsor the project in the subsequent year. On the other hand, the government’s focus on policy goals meant that initially some requests of governmental sponsors were disconnected from the needs and expertise of the maintainers. For example, initially the French government overlooked maintenance and asked the maintainers to work on DL, which was beyond their expertise. However, through an ongoing dialogue, these concerns were taken into consideration by the government’s AI strategy coordinators and the development priorities of the government were in turn adapted. Furthermore, maintainers appreciated the government’s requirement to secure additional public and private funding, which ensured the continued diversification of the funding base at a time when the French government was investing as much as €32 million into scikit-learn. As François remarked, “Overall, it’s good to have a panaché.”
5.2.2 Implications for companies
For companies, the findings are a timely reminder that by sponsoring a developer and/or funding a project, they can make a substantial difference to OSS maintainers, who struggle with limited resources and towering workloads [29]. As shown in Table 2, the scikit-learn project’s community includes several technology giants and industry leaders, such as Google, Meta, and AWS, which have not funded the project or sponsored maintainers to date. As one interviewee explained, while such companies contribute useful issues and pull requests, ultimately their contributions create more work for the already overstretched maintainers. It would be more helpful if companies such as these Big Technology companies sponsored a maintainer or funded the project by joining the consortium, which might be a low-effort intervention for a well-resourced industry giant with a high reward for the maintainers of scikit-learn.
5.2.3 Implications for governments
For governments, the findings underscore the critical importance of funding the maintenance of existing OSS and refraining from the novelty bias of exclusively funding the development of new OSS libraries or features.. This finding extends prior recommendations for the funding of scientific OSS [70]. Furthermore, it highlights the value of engaging in multi-stakeholder dialogue with OSS developers in order to inform the effective design and delivery of grants that strike a balance between the government’s policy goals and the daily realities of OSS developers. The scikit-learn maintainers explained that while they appreciated the recognition of the importance of scikit-learn in AI R&D in the French government’s AI strategy, they admitted they had faced difficulties in translating the government’s policy goals into practice. While they could understand the government’s rationale to strengthen digital sovereignty and national competitiveness, the government initially made demands that the maintainers viewd as either difficult to implement in practice or disconnected from their expertise, such as working on DL when their expertise was in ML. The maintainers suggested that this was in part due to a non-expert understanding of policymakers, which was remedied through ongoing multi-stakeholder dialogue between government officials, the scikit-learn maintainers, and other parties.
5.3 Threats to Validity
The validity of the study is evaluated by following guidance for case studies and qualitative research methods in the field of software engineering research [93, 76, 94].
5.3.1 Credibility
Credibility refers to the believability of qualitative research findings [94]. A number of steps were taken to enhance the credibility of the findings. First, multiple data sources—interviews, fieldwork, and secondary documents—were combined to triangulate findings and minimise the bias of individual sources [76]. Second, the prolonged engagement with the project’s maintainers over 17 months provided in-depth insights into the project’s evolving funding model, witnessing the ongoing negotiation with the government, as well as the arrival and departure of different consortium members. This longitudinal approach minimised the biases of snapshots at one moment in time. Third, findings were member-checked with research participants to enhance their accuracy [88]. Fourth, the community manager was invited to review a draft of the case study, further enhancing the accuracy and resonance of the findings [76]. However, data collection was constrained by data access challenges. A major obstacle was gaining access to interviewees, especially high-ranking stakeholders in the government and at sponsor companies, which is a common problem in “elite interviewing” [95]. For example, despite the generous support of the community manager and the offer of gift vouchers, only three representatives of commercial sponsors could be recruited for interviews.
5.3.2 Robustness
Robustness concerns the strength, reliability, and soundness of the study’s design, methods, and findings. A number of steps were taken to enhance the robustness. First, before the analysis, the theoretically-informed stakeholder groups were tested with interviewees to ensure their relevance, whilst accounting for nuances within and between stakeholder groups. This ensured that theoretical constructs were not imposed on the study that did not resonate with the research participants. Second, since the research data was collected over a prolonged time span of 17 months, the author took an iterative approach to the data analysis, combining an inductive, grounded theory analysis with a deductive approach based on the study’s theoretical framework throughout the period of data collection [87]. Third, the author maintained a social identity map throughout the research process to minimise the influence of personal biases on the study [89]
5.3.3 Transferability
Transferability concerns the generalisability of the findings. The single case study on scikit-learn threatens the generalisability of the findings. Certainly, it is easier to generalise from multiple case studies, which by design test direct replication [96] and mitigate the risk of case uniqueness [97]. For example, there are a number of unique characteristics of the scikit-learn project, such as the longstanding institutional support of Inria, the design of the consortium, and the specific support from the French government. However, these threats to the validity of this study can be tempered by considering that, contrary to statistical generalisation, case studies aim for analytical generalisation [76]. For example, there are aspects of the scikit-learn project that will resonate across OSS projects, including the maintainers’ experience of weaving the diverse interests of stakeholders into the project whilst seeking to safeguard the project’s community ethos, as well as the effective design and implementation of funding governance protocols. Furthermore, this study provides insights into the merits and challenges involved in the implementation of a public-private funding model in a community-led OSS project that may be useful for other projects considering similar approaches.
5.3.4 Dependability
Dependability refers to the consistency of the research process. To maximise the dependability of the research process, a systematic study protocol was employed for the collection, analysis, and storage of research data. Where consent was given, interviews were recorded and transcribed to aid analysis. When interview responses were ambiguous, follow-up discussions with the interviewees were initiated to obtain clarifications. In addition to interview data, detailed field notes from on-site visits and supplementary secondary documents facilitated data triangulation. Finally, the case study underwent thorough review, benefitting from the feedback from domain experts as well as peer review. The combination of these steps enhanced the dependability of the study.
6 Conclusion
Funding plays a critical role in supporting the sustainability of community-led OSS projects. This case study on scikit-learn makes two key contributions with implications for research and practice. First, the study contributes novel findings to the academic literature on emerging OSS funding models. It illustrates how the maintainers of scikit-learn have weaved public and private sources of funding into their project in the interest of ensuring the continued provision of scikit-learn as a digital public good, whilst employing governance protocols to safeguard the community ethos of the project. Second, it provides practical recommendations for OSS practitioners. For OSS developer communities, the case of scikit-learn shows that the diversification of funding can act as a useful mechanism for sustaining a community-led OSS project, whilst safeguarding the community ethos of the project and its independence from the influence of single funders. For companies, it is a timely reminder that by sponsoring an OSS developer or funding an OSS project they can make a significant difference to maintainers. For governments, it underscores the critical importance of refraining from the novelty bias of only funding new innovation and funding the maintenance of existing OSS. Overall, this study underscores the effectiveness of diversified funding models as a pragmatic approach to sustaining community-led OSS projects in competitive industries. Ultimately, the future of OSS funding depends on the development of sustainable strategies that align the interests of funders with the needs and norms of the global developer communities that build and maintain our open source digital infrastructure.
Acknowledgements
This study was supported by the UK’s Economic and Social Research Council [Grant number: ES/P000649/1].