### v4

# Public-private funding models in open source software development: A case study on scikit-learn

Abstract
Governments are increasingly funding open source software (OSS) development to support software security, digital sovereignty, and national competitiveness in science and innovation, amongst others. However, little is known about how OSS developers evaluate the relative benefits and drawbacks of emergent governmental funding for OSS. This paper explores this question through a case study on scikit-learn, a popular Python library for machine learning, which has been funded by public research grants, commercial sponsorship, micro-donations, and a €32 million grant announced in France’s artificial intelligence strategy. Through 25 interviews with scikit-learn’s maintainers and funders, this study makes two key contributions to research and practice. First, it contributes novel empirical findings on the effective design and implementation of a public-private funding model in an OSS project, as well as how the maintainers of scikit-learn have designed and employed governance protocols to balance the diverse interests of their funders and to safeguard their community ethos. Second, it offers practical lessons on funding in community-led OSS projects and makes recommendations to practitioners. The paper concludes with a discussion of the key recommendations.
Keywords Open source software, OSS funding models, machine learning, scikit-learn, OSS sustainability
1 Introduction
Open source software (OSS) has moved up the priority list of governments in light of concerns about digital sovereignty [1], software security [2], and national competitiveness in science and innovation [3], amongst others. The disclosure of the log4j vulnerability in November 2021 is widely seen as a turning point that mobilised national governments, from the United States of America (USA) to Germany, to mobilise public funds for the development and maintenance of OSS [4]. Increasing governmental involvement in funding OSS has been generally well-received by OSS developers. However, little is known about how they view the relative benefits and drawbacks of governmental funding compared to other types of funding. This paper explores this question through a case study on scikit-learn, a Python library for machine learning (ML), which is described as “the Swiss army knife of ML” due to its widespread use [5]. Its sustainability as a community-led project in the industry-dominated field of artificial intelligence (AI) [6] is partly attributed to its mixed funding model, including public research grants, commercial sponsorship, community donations, and the institutional backing from the French Institute for Research in Computer Science and Automation (Inria). In November 2021, in its national AI strategy, la “Stratégie IA”, the French government announced a €32 million grant to support the expansion of scikit-learn [7], marking a significant expansion of public funding for scikit-learn.
Through 25 interviews with scikit-learn’s maintainers and funders, this paper investigates the historical role of public and private funding in supporting the community-led OSS project. It explores the diverse interests of scikit-learn’s funders, and how they have aligned or conflicted with those of the maintainers (RQ1). The study also examines how the maintainers view the relative benefits and drawbacks of public and private funding (RQ2). Finally, it discusses practical lessons from scikit-learn’s funding model for OSS practitioners (RQ3). This paper makes two key contributions to research and practice. First, it provides novel empirical findings on the design and implementation of a public-private funding model in an OSS project. Furthermore, it sheds light on how the maintainers employed governance protocols to balance the diverse interests of their funders and to safeguard their community ethos. Second, it offers practical lessons on funding in community-led OSS projects. The paper concludes with a discussion of key recommendations to OSS developer communities, companies, and governments.
The paper has the following structure. First, it discusses prior work on the role of funding in supporting the sustainability of OSS projects. Second, it presents the study design and the case of scikit-learn. Then, it presents the results and discusses their implications for research and practice. The paper concludes with a discussion of the key recommendations to practitioners and suggestions for future research directions on OSS funding models.
2 Related Work
2.1 The Role of Funding in OSS Sustainability
The adoption of OSS is ubiquitous, estimated to be used in 96% of code bases [8] and constituting up to 90% of software stacks [9]. As a result, OSS is considered a critical component of the digital infrastructure that underpins modern society and the digital economy [10, 11, 12]. Given its importance, the health and sustainability of OSS projects is a concern for a diverse range of stakeholders. However, the health and sustainability of OSS projects is a complex and multi-faceted issue [13, 14, 15]. A range of factors are relevant, including but not limited to open governance and strong leadership [16, 17, 18, 19]; maintainer best practices, such as timely responses [20], mentoring contributors [21, 22], and recommending tasks [23]; a steady inflow of new contributors [24, 25, 22]; and diversity within the community [26]. Amongst these factors, the financial health of a project plays a crucial, but hitherto understudied, role in the sustainability of OSS projects, determining the extent to which maintenance is funded or reliant on volunteers [27].
However, money has long been a contentious subject amongst OSS developers. On the one hand, many developers take pride in the origins of the OSS movement as a social movement [28], describing it as a “programmers’ paradise” for “geeks”, “hackers”, and “hobbyists”, who do not contribute to OSS for money [29, 30]. Developers cite their political ideals [31], altruism [32], or their passion for spotting and quashing bugs in software [33] as intrinsic incentives for contributing to OSS. Furthermore, many developers complain that “money ruins everything” [34]. On the other hand, a plethora of OSS projects are unfunded and maintained by overstretched volunteers, who often struggle to keep up with the workload required to keep projects running and communities thriving [27, 35, 20].
The consequences of systemic issues of free-riding and under-investment in OSS are highlighted when developers spot major vulnerabilities that pose significant threats to digital systems in governments, critical infrastructure, and companies’ products and services. In particular, the discovery of vulnerabilities, such as the heartbleed bug in April 2014 or log4j vulnerability in December 2021, which was described as the “worst security problem in a generation” [36], shocked industry and governments alike. It highlighted the need for cross-sector stakeholders, including the public and private sectors, to fund the developer communities who build and maintain our open source digital infrastructure [4]. As Benjamin Bikinbine has argued, there is an urgent need “not just [for] investment in institutions, organisations, technologies, or innovations, but long-term and sustainable investment in the true source of their value: people” [37].
2.2 Private Funding for OSS Development
The private sector has been the largest funder of OSS to date, funding OSS developers and projects in both direct and indirect ways. The impact of commercial investments in OSS are gauged to be significant. According to one estimate, companies in the European Union (EU) invested around €1 billion in OSS in 2018, which generated between €65-95 billion for the EU’s GDP that year [38]. Direct funding methods include sponsorship of OSS foundations and consortia, which operate on a fee-paying membership model [39]. Foundations and consortia are known to play a critical role as “boundary organisations”, which through their open governance protocols facilitate collaboration between diverse volunteers and companies [16]. However, according to a survey of 369 OSS developers, only 12% of respondents reported that organisational funding for OSS via a foundation, a consortium, or an independent legal entity was either extremely or very useful to their development activity, while 5% reported it was ineffective [40].
Recently, companies have begun to establish FOSS Contributor Funds, which allow employees to nominate and vote on OSS projects that should receive funding [41], and represent a more “democratic” way for companies to fund the OSS projects that they use [42]. A shortcoming is that these funds typically award up to \$10,000 over a one-year time frame. Alyssa Wright from Bloomberg’s FOSS Contributor Fund [43] has acknowledged this shortcoming, arguing that “true sustainability comes from a multi-sided commitment” and “while financial support is important, real engagement through contribution and dialogue is as crucial to the ecosystem’s survival as writing a check”.
Other ways companies support OSS projects by sponsoring OSS developers or by allowing their employees to contribute to OSS projects at work [44, 45]. Sponsoring developers is a recognised strategy that companies employ to influence projects [46, 47] and to improve their reputation as OSS patrons, which in turn is considered to aid the recruitment of developers [48, 49, 50]. In the mid-2000s, it was estimated that 40% of OSS contributors were paid [51]. We should expect this number to be much higher today, with companies like Google, Microsoft, and Amazon accounting for a rapidly growing number of OSS developers and contributions to OSS projects [52]. In addition, companies invest in projects through donations and by joining project steering committees [53].
While OSS projects benefit from commercial participation, maintainers seek to manage the influence or the dominance of individual companies [54]. A popular strategy is to establish or transfer projects to foundations [39, 55], which through open governance protocols limit the dominance of single companies [56] and facilitate collaboration between diverse contributors [16, 57]. Crucially, these organisations enable OSS developers to “negotiate their relationship with companies and, when necessary, to defend their commons-based resources from unwanted influence” [37].
2.3 Micro-Donations for OSS Development
Individuals or organisations support OSS developers and project by making micro-donations to individual developers and OSS projects. For instance, the GitHub Sponsors scheme, launched in 2019, allows users to make one-off or recurring payments to individuals or organisations with sponsored profiles [58]. According to prior work, common incentives for donations via GitHub Sponsors include expressing appreciation, recognising a developer’s work, and encouraging future contributions [30]. By comparison, common incentives for maintainers to accept donations are to receive recognition for their work or to provide a side-income; while those who do not participate in GitHub Sponsors state that they do not need to be sponsored or that they do not contribute to OSS for money [30]. Furthermore, tther work shows that the amount that individuals donate is associated with the length of their use of an OSS [59]. The impact of micro-donations is still contested. On the one hand, they have been shown to shorten response times to issues [60, 61] and to increase maintenance-related activities [62]. On the other hand, overall the impact appears trivial [63] and donations tend to be ad-hoc, small, or legally complicated for individuals to receive [64].
2.4 Public Funding for OSS Development
The public sector has indirectly or directly funded OSS since the origins of the World Wide Web [65] through research grants for (scientific) OSS [66, 67, 68] as well as bespoke OSS funding bodies, such as the Open Technology Fund in the USA, the Next Generation Initiative at the European Commission, or the Sovereign Tech Fund (STF) in Germany [4]. Furthermore, OSS has recently moved up the priority list of governments in light of concerns about the security of OSS [69, 2]; its role as critical digital infrastructure [10, 11, 12]; and its importance for digital sovereignty [1, 70] and national competitiveness in science and innovation [3]. In Europe, the concern for digital sovereignty is particularly salient. For example, the European Commission’s OSS strategy argues that OSS can give “Europe a chance to create and maintain its own, independent digital approach and stay in control of its processes” [71].
Public sector support for OSS is also a recognised strategy to stimulate competition in domestic software markets, which might otherwise be monopolised by industry giants [72]. For example, it has been estimated that the Circulaire 5608, a French law requiring government agencies to favour OSS when procuring software, led to an increase of nearly 600,000 OSS contributions from French developers per year as well as yearly increases in the national competitiveness of the French IT market [3]. Beyond governmental adoption of OSS, government investments in OSS are often targeted at stimulating national competitiveness. For example, in June 2023, President Emmanuel Macron of France announced a €40 million fund to create a “digital commons” to support the development of open source large language models to support domestic startups and to challenge the dominance of Big Technology companies in the AI industry [73].
The discovery of the log4j vulnerability in November 2021 marked a turning point for governmental interest and involvement in funding OSS development. In light of the magnitude of the security concerns [69], governments rapidly mobilised to address the threats posed by OSS vulnerabilities to their critical infrastructure and digital economies. For example, in January 2022, the White House convened cross-sector stakeholders, including government agencies, Big Technology companies, and OSS foundations, to discuss the urgent need to improve the security of OSS, recognising its indispensable role as digital infrastructure whose maintenance largely depends on volunteers [74]. Across the pond, in June 2022, the newly formed European Working Team on the Digital Commons, comprising representatives from 19 member states of the EU and the European Commission, published its inaugural report, highlighting the importance of funding OSS for bolstering the digital sovereignty of Europe [75]. In October 2022, Germany established the STF, allocating a budget of €11.5 million for its first year alone to fund OSS maintenance in the interest of digital sovereignty, software security, innovation, and digital democracy [76]. These developments underscore the growing recognition amongst policymakers of the public value of funding OSS development and maintenance [4].
While there is a growing recognition of the need for sustainable funding for OSS by governments, little is known about how OSS developers evaluate the relative merits and drawbacks of governmental involvement in funding OSS development compared to other types of funding that have primarily supported OSS development so far, or the impact thereof on OSS development and maintenance. This study addresses this gap by examining the case of scikit-learn and how its maintainers view the benefits and drawbacks of its public and private funders. By focusing on the perspectives of OSS developers, this study contributes practical insights on this timely question in the debate about OSS sustainability.
3 Study Design
3.1 Research aims
This paper investigates the emerging role that governments are playing in funding OSS development and maintenance, and how OSS developers view the relative merits and drawbacks of governmental funding through a case study on the scikit-learn project. In software engineering research, a case study “is an empirical enquiry that draws on multiple sources of evidence to investigate one instance (or a small number of instances) of a contemporary software engineering phenomenon within its real-life context” [77]. Single case studies are suitable when one seeks to provide in-depth insights, generate new hypotheses, or develop theories about a critical, extreme, unusual, or revelatory phenomenon [78]. The scikit-learn project was selected as an exploratory case that could provide in-depth insights into the role of the French government in supporting this OSS project and how its maintainers view the relative merits and drawbacks of the government’s funding compared to the other funding sources that have supported the project to date.
To this end, the study focuses on the following research questions (RQ):
  • RQ1: What are the interests of the public and private funders of scikit-learn, and how have they aligned or conflicted with the interests of the scikit-learn maintainers?
  • RQ2: How do the maintainers of scikit-learn view the relative benefits and drawbacks of public and private funding for their OSS project?
  • RQ3: What practical lessons can be drawn from scikit-learn’s funding model for OSS practitioners?
3.2 Case presentation
scikit-learn is a Python library that implements ML algorithms for classification, regression, and clustering, as well as tools for data pre-processing, model evaluations, and data visualisation. Initially it was started as scikits.learn by David Courapeau during a Google Summer of Code project in 2007. After a dormant period, it was relaunched as scikit-learn by Fabian Pedregosa, Gaël Varoquaux, Alexandre Gramfort, and Vincent Michel with the backing of INRIA, with the first public release in February 2010. It has since become one of the most popular OSS libraries for ML [INRIA_2019_2020], described as “the Swiss army knife of ML” due to its widespread use [INRIA_2019_2020]. For example, in November 2023, it had 2,752 contributors and 61,085 GitHub stars [79]. The success and sustainability of this community-led OSS project is noteworthy, considering the extent of industry dominance in ML/AI research [6] and OSS ecosystems [80, 81].
scikit-learn’s sustainability is partly credited to its funding model, which has combined research grants, private sponsorships via the scikit-learn consortium, and community donations [82, 83]. The consortium was established under the Inria Foundation in 2018 in order to secure employment for maintainers, to sustain the quality of software, and to develop new features [83]. It comprises companies that pay annual membership fees. Since its founding, it has included Dataiku, Microsoft, Nvidia, Intel, AXA, Boston Consulting Group, BNP Paribas Cardif, Hugging Face, Nvidia, Fujitsu, and Chanel. Other companies, such as Quansight Labs, have indirectly funded the project by sponsoring maintainers. The project also benefits from donations from community members via NUMFOCUS and student projects have been sponsored by the Google Summer of Code programme. In November 2021, the French AI strategy, la “Stratégie IA”, announced a €32 million grant to support scikit-learn’s development [7]. This significant public investment alongside scikit-learn’s existing funding sources makes it a suitable case for the research aims of this exploratory case study.
3.3 Data & analysis
3.3.1 Data collection
To answer the research questions, this study employs a mixed-methods approach, comprising document analysis and 25 semi-structured interviews with scikit-learn maintainers, commercial sponsors, and government officials between November 2021 and April 2023. Most interviews took place during on-site visits to the scikit-learn team’s office in Paris. The sampling approach sought to ensure complete coverage of the various stakeholder groups in the project—that is, the OSS developers (i.e. the scikit-learn maintainers), commercial sponsors (i.e. the consortium members), and public funders (i.e. Inria and the French government)—as well as diversity within each group. Interviews were conducted with maintainers with various tenures in the project, including some who had been involved for over 10 years and recent university graduates. One maintainer worked for a company that was a consortium member but spoke on his own behalf. From the consortium, interviews were conducted with three representatives from three companies with different sponsorship tiers, organisational sizes, and different sectors. In addition, interviews were conducted with two government officials who were responsible for the coordination of the AI strategy. The interviews were semi-structured, combining general questions about scikit-learn’s funding model and tailored questions for each interviewee [84]. To inform tailored questions for each maintainer, their contribution histories on GitHub’s Contributor Insights Page were reviewed; and for the funders, public information about their involvement in funding scikit-learn was reviewed. The author conducted and recorded every interview to aid the analysis and enhance the validity of the findings [78].
The interview data was supplemented and triangulated through detailed field notes from on-site visits and secondary documents to strengthen the validity of the findings [78]. The field notes stemmed from two on-site visits to the scikit-learn office in Paris-Saclay in March 2022 and 2023; two on-site visits to the offices of sponsors in September and December 2022; and attendance at the Global Partnership on AI Summit in November 2021, which was hosted by the French government and attended by a co-founder of scikit-learn. The secondary documents included both publicly available and non-disclosed documents, which interviewees shared concerning the French government’s AI Strategy and its funding package for scikit-learn. Additional information about scikit-learn and its funding model was collected from internet searches, included the scikit-learn website and company blogs, or provided by interviewees.
3.3.2 Qualitative data analysis
Throughout the 17 months of data collection, an iterative, integrated approach was employed to identify common themes in the interview data, field notes, and secondary documents [85]. Given the exploratory aims of this study, the integrated approach combined the application of grounded theory analysis [86] and a deductive, organising framework informed by prior literature about the different interests and involvement models of individuals, companies, and governments in the development of OSS [37, 48, 72, 16]. Thematic codes were continuously developed as data was collected, and in turn translated into themes until the point of saturation [87]. The resulting themes were member-checked to ensure their accuracy and resonance with research participants [88]. Throughout the research process, a social identity map was maintained by the author to exercise reflexivity regarding the potential influence of personal biases stemming from the author’s views and relationships built with the scikit-learn team during the 17 months of data collection [89].
It should be noted that each stakeholder group was not assumed to be homogenous; instead, it was accepted that there could be divergent or even conflicting interests within each stakeholder group. For instance, the scikit-learn consortium comprises a range of companies, including startups and multinational corporations from different sectors. Prior work indicates that companies primarily pursue strategic interests by participating in OSS [90], and it was not assumed that the consortium members shared the same strategies in funding scikit-learn. Similarly, public funders, including Inria and government officials who were responsible for France’s AI strategy, were grouped together solely by their belonging to the French public sector, despite obvious differences between these stakeholders. Even the maintainers were a heterogeneous group, combining employees of Inria and external companies, each with their unique paths to and histories in the scikit-learn project. These theoretical groupings were used solely for analytical purposes and were iteratively refined during the 17 months of data collection to ensure their coherence and suitability for the analysis.
4 Results
The results are organised into three parts, sequentially responding to RQ1 on how the interests of scikit-learn have aligned with or differed from those of the maintainers, and RQ2 on the relative benefits and drawbacks of public and private funding from the perspective of the scikit-learn maintainers. Finally, practical lessons and recommendations for practitioners based on the funding model of scikit-learn RQ3 are discussed in the subsequent Discussion section.
4.1 Stewards of the Community: The Roles & Interests of the scikit-learn Maintainers
scikit-learn’s maintainers are a diverse group employed by INRIA, public universities, and software companies. While some members of the core team work on scikit-learn full-time, others balance their roles as maintainers with their full-time jobs. Several maintainers also held academic positions at universities and research labs. The maintainers explained they perform a range of technical and social tasks, including reviewing code, mentoring contributors, and community building. François Goupil, the community manager, explained, “Of course, software development itself is the major part of the workload—going through issues, pull requests, and responding to community feedback—but lots of other work goes into the project. Community animation is a huge part. After all, it’s a community project”. In addition, sponsored developers from companies make focused contributions, such as the development of new features, which are agreed upon by the maintainers and the sponsoring companies. For example, at the time of this research, Nvidia sponsord a full-time maintainer who was building a plug-in to enable alternative computational backends for scikit-learn to enable compatibility with Nvidia’s GPU AI hardware accelerators.
Several maintainers expressed a strong sense of stewardship for the project, whose strength they viewed as rooted in its diverse community, comprising contributors from across the world (see Table 1) and companies (see Table 2). Julien Jerphanion, a maintainer, highlighted this community ethos, noting that, “the major part of the labour [in the project] is based on benevolence of people working in their free time and not asking to get paid”. Furthermore, Gaël Varoquaux, a co-founder and maintainer, explained that the community was at the heart of the project’s DNA and that major decisions about the future of the project remained community-driven despite their diverse funders. Similarly, Adrin Jalali, a maintainer, commented, “Even as a [maintainer], I don’t have enough credit to just be like, ‘Hey, we should do this.’ It’s a slow project with hundreds of 1000s of users. You can’t just change things”.
Table 1: Most active countries in the scikit-learn community [79]
Stargazer Issue Creator Pull Request Creator
1 China (23.7%) USA (33.7%) USA (34.0%)
2 USA (21.8%) Germany (7.5%) India (8.9%)
3 India (7.9%) UK (7.0%) Germany (8.5%)
4 Germany (4.4%) India (6.8%) France (6.9%)
5 Brazil (3.8%) France (5.5%) UK (6.4%)
6 UK (3.7%) China (5.4%) Canada (4.5%)
7 Canada (3.3%) Canada (4.5%) China (2.6%)
8 France (3.2%) Netherlands (2.5%) Japan (2.5%)
9 Japan (2.5%) Switzerland (2.2%) Switzerland (2.3%)
10 Russia (1.3%) Australia (2.0%) Netherlands (2.2%)
Table 2: Most active countries in the scikit-learn community [79]
Stargazer Issue Creator Pull Request Creator
1 Google (1.5%) Google (2.5%) Google (3.7%)
2 Microsoft (1.5%) Microsoft (2.0%) Microsoft (2.5%)
3 Tencent (0.9%) Amazon (0.7%) INRIA (0.9%)
4 Alibaba (0.8%) DeepMind (0.7%) Meta (0.8%)
5 Amazon (0.8%) IBM (0.6%) Amazon (0.8%)
6 ByteDance (0.8%) Johns Hopkins (0.6%) Dataiku (0.7%)
7 Baidu (0.6%) Tencent (0.5%) UC Berkeley (0.6%)
8 Meta (0.5%) ETH Zurich (0.5%) Shopify (0.6%)
9 AWS (0.4%) Uni. of Washington (0.4%) MIT (0.6%)
10 Tsinghua University (0.4%) NeuroData (0.4%) AWS (0.6%)
In addition, the maintainers expressed their personal interests in maintaining scikit-learn. Several maintainers explained they became maintainers because of their scientific interest in ML, and they liked the scientific culture of the project. Many of the maintainers had academic backgrounds, some still worked in research labs, and some still taught at universities. For example, Julien Jerphanion said that he enjoyed working at the intersection of statistical theory and software engineering, which was lacking in his previous job in industry. In addition, an anonymous maintainer explained that while he may not become a millionaire, he had a comfortable life and enjoyed the challenge of the job. For many, being a maintainer of scikit-learn was also a source of pride, given the project’s reputation.
Beyond maintaining scikit-learn, several maintainers mentioned the importance for them to contribute to the wider scientific Python ecosystem and to support their dependencies, such as numpy and scipy. During my on-site visit in September 2022, the maintainers had just returned from EuroSciPy, an annual gathering of the European scientific Python community. One anonymous maintainer spoke fondly of “putting faces to GitHub IDs”. François Goupil explained that beyond collaborating on the code base, they collaborate with other projects in other ways, such as by applying for grants together. In addition, some of the maintainers had initiated “satellite projects”, such as joblib and skrub. However, their ability to support the ecosystem is constrained by their limited capacity. Julien Jerphanion explained that while projects like scikit-learn, numpy, and scipy are very important in ML/AI, “There are many which are not visible to the public, for example OpenBLAS which is fundamental and only maintained by two people, who have very specific expertise. This is tremendous software work, but it is invisible to the public. Projects like this are cornerstones for many projects. They are the foundations; they are, like, infrastructure — you do not see them if there is no problem in your daily life”.
4.2 Public Funding: From Research Grants at INRIA to “la Stratégie IA”
Public funding has been a cornerstone of scikit-learn’s development since its beginning. Starting as a Google Summer of Code project in 2007, the library was later revived by researchers at INRIA. Over the past 13 years, INRIA had provided an estimated €1.5 million in support through staff salaries, research grants, office space, computing resources, and event sponsorship. As Olivier Grisel, a maintainer, noted, “[Public funding] is not new with the AI strategy”. However, the announcement of a €32 million grant in France’s AI strategy, la “Stratégie IA”, supporting the expansion of scikit-learn and OSS tools for data science, marked a significant expansion of public funding for scikit-learn. A government official explained that the grant seeks to facilitate AI adoption across the economy and to enhance French competitiveness in AI R&D, while also supporting the “digital sovereignty” of France and Europe.
The maintainers were delighted to receive the funding and recognition from the government. Adrin Jalali highlighted the stability it provided for the project, enabling longer-term planning around hiring and the technical roadmap. Julien Jerphanion characterised it as “a good signal that the French state is getting involved in funding projects like this”. François Goupil particularly commended the requirement to seek matching funds from both public and private sources across the European Union, viewing this as an important safeguard for the project’s independence: “I think it would be dangerous for us to be exclusively funded by the private sector or to be exclusively funded by the French government, because we have many good contributors who are not French”.
At the same time, challenges have emerged in aligning broad policy goals with the day-to-day realities and expertise of the maintainers. Initial requests from officials to expand scikit-learn into deep learning to provide a “sovereign” alternative to PyTorch and TensorFlow were resisted, with maintainers insisting on focusing on the project’s core strength in classical ML and the need to allocate sufficient funding and capacity to maintenance. The government officials ultimately deferred to their expertise, with one official acknowledging, “We understand that scikit-learn is specialised in ML, and it should not compete against TensorFlow and PyTorch”.
However, other tensions have proven thornier. Requirements to simultaneously develop OSS tools for data science whilst not competing with domestic companies have left maintainers in a bind, as the resulting tools inherently provide open source alternatives to the proprietary offerings of companies. Maintainers have also expressed frustration with the slow pace of disbursement, with funds still not arriving 17 months after the announcement by the time of the last interview. While scikit-learn’s diverse funding model provided a buffer, they worried that such sluggishness could threaten the sustainability of more financially precarious OSS projects.
Another official explained that the funding decision was informed by a multi-stakeholder process involving experts from research and industry. “It was not just the government. It was a dialogue between several actors—with the AI strategy team, the ‘Confiance IA’ project,111“Confiance IA” is a government-sponsored programme that aims to develop and industrialise trustworthy AI technologies in France. INRIA, with scikit-learn, and so on”. He confessed they had faced hurdles within the government: “Financing open source means financing a product that cannot be valued on the market and a priori it is not clear what its success will be. So, if [we want] the government to finance it in a substantial way—here we are talking about €32 million—the government wants to know where the money is going and to achieve a measurable return on investment”. Ultimately, it took months of coordination and negotiation to get the grant approved by the finance ministry.
Two consortium members raised concerns that such a substantial government grant could politicise scikit-learn, potentially influencing it to prioritise French policy objectives over the interests of the community. One anonymous representative underlined the need to balance the French government’s concern about digital sovereignty with the norms of the global community, while another argued that imposing French policy priorities on the project would not be received well by the community. These concerns were acknowledged by the maintainers, who insisted that the community ethos of the project remained at the heart of their discussions with the government.
4.3 Private Sponsorship: The Role of the scikit-learn Consortium
scikit-learn has received significant support from the private sector through the scikit-learn consortium and sponsorship of maintainers by companies like Nvidia, HuggingFace, and QuantSight. By creating the consortium, François Goupil explained they had sought to provide a more stable source of funding beyond research grants and to hire full-time maintainers, which was not possible previously.
Companies in the consortium have pursued a range of social, economic, and technological goals. By joining the consortium, companies are able to support a software tool that they use, whilst building goodwill and a reputation as OSS advocates, which an important signal for hiring in a competitive labour market. François Goupil explained, “A lot of people come from this open source background and if you see that a company is promoting such an open source project, you might say, ’Okay, maybe, I would be happy working with them. They could be okay with me contributing to OSS at work”. He viewed this incentive as more important for companies that did not already have a strong reputation in ML/AI field, since it would make them more credible and attractive to job-seekers.
For their part, consortium members cited a mix of pragmatic and principled reasons for funding scikit-learn. On a basic level, it allows them to support the maintenance of a tool they use in their products and services. Mike McCarty from Nvidia shared that his concern about unmaintained dependencies “keeps [him] up at night”, citing numpy as a critical library that had been under-maintained for some time. Leo Dreyfus-Schmidt from Dataiku expressed a similar sentiment, who stated, “We are very much aware that we rely on open tools like scikit-learn” and highlighted the importance of “giv[ing] back to the community”. An anonymous consortium member was even more direct: “Companies pay millions in commercial licenses, so why not support the OSS projects we use¿‘
However, discussions with maintainers and consortium members underlined that consortium membership is not simply a charitable contribution. In exchange for their sponsorship, companies gain a voice in scikit-learn’s development through bodies like the Technical Committee and Advisory Board, which are venues for surfacing industry use cases and needs. Leo Dreyfus-Schmidt characterised this as a “win-win”, allowing scikit-learn to benefit from “the perspective of people in industry, which…is a very different type of feedback from the one they get from the open source community”, whilst giving commercial members direct input into the project.
In practice, this dynamic has yielded a mix of synergies and tensions. In some cases, companies’ business interests are compatible with scikit-learn’s technical roadmap. For example, in addition to its consortium membership, Nvidia has sponsored a maintainer to build a plug-in that enables alternative computational backends for scikit-learn. Mike McCarty explained, “One thing that Nvidia is very interested in is scikit-learn’s effort to accelerate with GPUs. I would love to see at the end of the day scikit-learn to be accelerated natively with a very nice user experience, kind of like what we have with PyTorch, which is the canonical example of a library that you can install and if you happen to have a GPU, you can accelerate it without changing your code”. Guillaume Lemaitre, a maintainer, described Nvidia’s simultaneous sponsorship of a maintainer and consortium membership as “a win-win situation”.
In other cases, however, sponsor requests to shape the project’s direction or burnish their own reputations have run afoul of project norms. Two anonymous maintainers explained that in the past, a sponsor had requested modifications to the technical documentation that amounted, in their view, to free advertising for the company. This self-promotional request, they explained, was not acceptable, especially not in the technical documentation of the project. While disputes of this kind have been rare, one anonymous maintainer mentioned that disputes had led to the termination of sponsorship by one company in the past. To manage the interests and demands of consortium members, the scikit-learn maintainers employ carefully designed governance protocols that limit funders’ influence and prioritise community control. The Technical Committee, for instance, is barred from overriding decisions that achieve rough consensus amongst the maintainers. Adrin Jalali explained, “We don’t want the technical committee to have power over the core contributor team as long as [the maintainers] have a consensus. If there’s two thirds majority in cast vote, for whatever vote we do, then the technical committee doesn’t step in…because the power is given to the [maintainers] and that’s by design”.
These governance measures are important to maintain the trust of the community and to mitigate fears that the project had been bought out. François Goupil explained that as a community-led OSS project, they communicate to consortium members that their feedback is valued but they cannot guarantee that they will implement it. For example, the maintainers organised a workshop on MLOps because this was a topic that the consortium members had expressed an interest in. “MLOps will probably end up in the technical roadmap in some way…but it’s a discussion and in the end the team will decide if it’s worth it or not”, he explained. Through these governance protocols, the maintainers ensure that the project remains community-driven while still considering and incorporating feedback from the consortium. As François Goupil put it, “We are a community with different interests”, and preserving that pluralism is key to the project’s long-term sustainability and independence.
Finally, even if relationships with their sponsors can be troublesome at times, the maintainers were grateful for their support and for not free-riding on scikit-learn, which many companies do. During an on-site visit, François Goupil showed me a dashboard, comprising world maps of contributors and charts of contributions from companies. Upon observing that the most active organisations were Chinese and American companies (see Table 2), I asked if these insights had influenced his opinion about whether they contributed sufficiently. He responded frankly: while such companies contribute useful issues and pull requests, their contributions ultimately create more work for the maintainers. He explained that it would be more helpful if these companies sponsored a maintainer or joined the consortium in order to provide more capacity to the maintainers to process the extra workload created by these contributions.
5 Discussion
5.1 Implications for Research
5.1.1 The design and role of public-private OSS funding models
This study contributes novel findings on the roles of public and private funding in OSS development and maintenance. It reveals the diverse interests of funders, from the economic and technological goals of companies [48] to the policy objectives like digital sovereignty [1], national competitive in science and innovation [3], and economic growth [72] for governments. While the various stakeholders pursue different goals, they are not necessarily in conflict, as they share the goal of sustaining scikit-learn. For example, while the maintainers and the companies do not necessarily share the government’s policy goals or digital sovereignty concerns, its long-term investment in scikit-learn aligns with the maintainers’ and companies’ interest in its sustainability for research and innovation purposes. Similarly, while the maintainers and the government officials do not necessarily share the companies’ business interests in scikit-learn, the project benefits from companies’ feedback and funding for development and maintenance. The study also highlights the role of the governance mechanisms to balance the interests of the community and the companies in the consortium, extending prior work on the role of foundations as “boundary organisations” that mitigate the dominance of companies and enable collaboration between OSS developers and companies [16, 54].
5.1.2 Research design for future research on OSS funding
There is still a dearth of research on OSS funding, relative to other topics on the economics of OSS [91]. With the emergence of new funding models, from FOSS Contributor Funds to governmental funds like Germany’s STF, it is timely to empirically investigate OSS funding from multiple angles, including funders, incentives, funding models, and impact(s), amongst others. To this end, this paper contributes novel insights and a research design that can inform future studies on OSS funding. As discussed below, while the uniqueness of scikit-learn limits the generalisability of the case-specific findings, there are aspects of the scikit-learn experience that will resonate with OSS developers in general, such as the processes of balancing the varied interests of funders while preserving community values.
5.2 Implications for Practice
5.2.1 Implications for OSS Developer Communities
The case of scikit-learn shows that a diversified funding model can be an effective strategy for safeguarding a project’s independence and community-driven ethos whilst benefiting from both public and private support. Furthermore, public and private funding come with their own challenges. On the one hand, commercial sponsors often focus on specific technical issues related to maintenance and feature development, and their demands may sometimes conflict with project goals or norms. On the other hand, government funders may prioritise high-level policy objectives over the activities or expertise of the project. Navigating these tensions requires ongoing dialogue and negotiation to align goals and ensure a sustainable balance of funding sources. Ultimately, maintainers must be prepared to advocate for and defend their project’s needs and values. As the scikit-learn case demonstrates, a diverse funding portfolio can provide both stability and leverage in these discussions.
5.2.2 Implications for Companies
For companies, the findings serve as a timely reminder that sponsoring OSS maintainers or projects can make a substantial difference to maintainers, who often struggle with limited capacity and towering workloads [27]. As shown in Table 2, several major technology companies, such as Google, Meta, and AWS, have benefited significantly from scikit-learn but have not yet directly funded the project or its maintainers. While these companies may contribute useful issues and pull requests, their engagement can also create additional work for already overstretched maintainers. Sponsoring a maintainer or joining a project’s funding consortium, even at a relatively low level, could be a high-impact way for well-resourced companies to support the OSS projects that they use and depend on. By stepping up to provide more direct financial support, companies can help ensure the sustainability and health of critical OSS infrastructure.
5.2.3 Implications for Governments
For governments, the findings highlight the importance of funding the maintenance of existing, widely-used OSS infrastructure and not just the development of new libraries or features, extending prior recommendations for the funding of scientific OSS [68]. To ensure that funding is aligned with project needs, governments should engage in multi-stakeholder dialogue with OSS developers. This consultation is crucial for designing grants that balance policy goals with the expertise of maintainers. In the case of scikit-learn, initial government requests around expanding into deep learning were misaligned with the project’s scope and capabilities. Ongoing conversations were necessary to bridge this gap and ensure that funding supported both overarching policy objectives and essential project functions. Governments should proactively seek input from the OSS community to avoid such misalignments and to craft funding programs that effectively support the sustainability of critical (scientific) OSS projects.
5.3 Threats to Validity
The validity of the study is evaluated by following guidance for case studies and qualitative research methods in the field of software engineering research [92, 78, 93].
5.3.1 Credibility
Credibility refers to the believability of qualitative research findings [93]. A number of steps were taken to enhance the credibility of the findings. First, multiple data sources—interviews, fieldwork, and secondary documents—were combined to triangulate findings and minimise the bias of individual sources [78]. Second, the prolonged engagement with the project’s maintainers over 17 months provided in-depth insights into the project’s evolving funding model, witnessing the ongoing negotiation with the government, as well as the arrival and departure of different consortium members. This longitudinal approach minimised the biases of snapshots at one moment in time. Third, findings were member-checked with research participants to enhance their accuracy [88]. Fourth, the community manager was invited to review a draft of the case study, further enhancing the accuracy and resonance of the findings [78]. However, data collection was constrained by data access challenges. A major obstacle was gaining access to interviewees, especially high-ranking stakeholders in the government and at sponsor companies, which is a common problem in “elite interviewing” [94]. For example, despite the generous support of the community manager and the offer of gift vouchers, only three representatives of commercial sponsors could be recruited for interviews.
5.3.2 Robustness
Robustness concerns the strength, reliability, and soundness of the study’s design, methods, and findings. A number of steps were taken to enhance the robustness. First, before the analysis, the theoretically-informed stakeholder groups were tested with interviewees to ensure their relevance, whilst accounting for nuances within and between stakeholder groups. This ensured that theoretical constructs were not imposed on the study that did not resonate with the research participants. Second, since the research data was collected over a prolonged time span of 17 months, the author took an iterative approach to the data analysis, combining an inductive, grounded theory analysis with a deductive approach based on the study’s theoretical framework throughout the period of data collection [87]. Third, the author maintained a social identity map throughout the research process to minimise the influence of personal biases on the study [89]
5.3.3 Transferability
Transferability concerns the generalisability of the findings. The single case study on scikit-learn threatens the generalisability of the findings. Certainly, it is easier to generalise from multiple case studies, which by design test direct replication [95] and mitigate the risk of case uniqueness [96]. For example, there are a number of unique characteristics of the scikit-learn project, such as the longstanding institutional support of Inria, the design of the consortium, and the specific support from the French government. However, these threats to the validity of this study can be tempered by considering that, contrary to statistical generalisation, case studies aim for analytical generalisation [78]. For example, there are aspects of the scikit-learn project that will resonate across OSS projects, including the maintainers’ experience of weaving the diverse interests of stakeholders into the project whilst seeking to safeguard the project’s community ethos, as well as the effective design and implementation of funding governance protocols. Furthermore, this study provides insights into the merits and challenges involved in the implementation of a public-private funding model in a community-led OSS project that may be useful for other projects considering similar approaches.
5.3.4 Dependability
Dependability refers to the consistency of the research process. To maximise the dependability of the research process, a systematic study protocol was employed for the collection, analysis, and storage of research data. Where consent was given, interviews were recorded and transcribed to aid analysis. When interview responses were ambiguous, follow-up discussions with the interviewees were initiated to obtain clarifications. In addition to interview data, detailed field notes from on-site visits and supplementary secondary documents facilitated data triangulation. Finally, the case study underwent thorough review, benefiting from the feedback from domain experts as well as peer review. The combination of these steps enhanced the dependability of the study.
6 Conclusion
This study on the funding model of scikit-learn makes two key contributions to research and practice. First, it contributes novel findings on emerging OSS funding models, illustrating how the maintainers of an impactful community-led OSS project have combined public and private funding to sustain the development and maintenance of the project, whilst safeguarding its community ethos. Second, it offers practical lessons for the funding of community-led OSS projects, and it makes recommendations to various OSS practitioners. In particular, it highlights effectiveness of diversified funding models as a method to sustain community-led OSS projects in industry-dominated R&D fields, such as AI.
Acknowledgements
This study was funded by the UK’s Economic and Social Research Council [Grant number: ES/P000649/1]. The author thanks the scikit-learn maintainers and funders for participating in this study.